OpenStudy (anonymous):

all solutions to 3 sec x - 2 = 0

OpenStudy (zepp):

\[\large \begin{align} 3\sec(x)-2 &= 0\\ 3\sec(x)&=2\\\sec(x)&=\frac{2}{3} \end{align}\]

OpenStudy (zepp):

\(\large \sec(x)=\frac{1}{\cos(x)}\)

OpenStudy (zepp):

\[\large \frac{1}{\cos(x)}=\frac{2}{3}\\\large \cos(x)=\frac{1}{\frac{2}{3}}=1(\frac{3}{2})=\frac{3}{2}\]

OpenStudy (zepp):

\(\large x = \arccos\frac{3}{2}\)

OpenStudy (anonymous):

@_oftheSOUL you like some big black mushroom tip? ;P